OpenStudy (johnnydicamillo):

limit as x approaches infinity of (x)/log_2(3^x))

sam (.sam.):

\[ \lim_{x\to \infty} \frac{x}{\log _2\left(3^x\right)}\]

OpenStudy (johnnydicamillo):

right

OpenStudy (johnnydicamillo):

I am trying to compare asymptotic behavior

OpenStudy (johnnydicamillo):

so I know this is an infinity over infinity case, how would I find the derivative of the denominator?

OpenStudy (agent0smith):

You can either use the "rules" for log deriv or say to hell with the rules! Actually no, use the other rules... \[\large \log a^x = x \log a\]easy now huh ;)

OpenStudy (agent0smith):

\[\lim_{x\to \infty} \frac{x}{\log _2\left(3^x\right)} = \lim_{x\to \infty} \frac{x}{x\log _2\left(3\right)}\]

OpenStudy (johnnydicamillo):

\[\frac{ 1 }{ 3 \ln(2) }\]

OpenStudy (agent0smith):

https://cdn.meme.am/instances/500x/55941579.jpg